The price of oil is $30 per barrel and the price elasticity is constant and equal to -0.5. An oil embargo reduces the quantity available by 20 percent. Use the arc elasticity formula to calculate the percentage increase in the price of oil.


1
Expert's answer
2021-02-03T17:45:51-0500

The price elasticity of demand using the arc elasticity formula can be found as follows:


"E_d^{arc}=\dfrac{\%\ Change\ in\ Quantity\ Demanded}{\%\ Change\ in\ Price},""-0.5=\dfrac{-0.2}{\%\ Change\ in\ Price},""\%\ Change\ in\ Price=\dfrac{-0.2}{-0.5}\times100\%=40\%."

Therefore, the percentage increase in price is 40%.